Shift Summary for Day Shift, 6/25/99

S. Elizabeth Turner (eturner@kudzu)
Fri, 25 Jun 1999 15:32:03 -0500

Shift Summary for Day Shift, 25/Jun/99

On Shift:
SET,AA

Hours of beam delivered at all intensities: 6
Hours of beam delivered at nominal intensity: 6
Hours of E832 data (special runs not included): 6
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
Beam down ~2.5 hours due to TeV spill problems.

========================================================================

Runs ended this shift:
13724: BEAM_832_TEST (no E-W-U-D logic in charged trigger)
13725: BEAM_832 (beam studies)
13726: BEAM_832 (beam studies)
13727: BEAM_832 (beam studies)
13728: BEAM_832 (beam studies)
13729: BEAM_832 (beam studies)
13730: BEAM_832 (beam studies) -- in progress at end of shift

Tapes brought to FCC this shift:
(including tapes for online split)
None
========================================================================
Accesses during shift:
None
========================================================================
Overview of shift:
Poor quality beam with several downtimes during first half of shift.
Rick C. does beam studies for second half with good beam.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

13724 8:15 no beam
58 8:40 beam returns
9:00 no beam
9:41 AA picks up 2365 module from PREP on orders from
people looking at V1E problem
179 11:10 beam returns
12:30 end run

13725 12:30 begin run, Rick C. commences beam studies

13726 13:17 beam studies

13727 beam studies

13728 14:15 beam studies
14:20 SET changes 2 REG HVs (-3V) before start of run

13729 15:00 beam studies

13730 15:10 beam studies

All day:
DC distance track-hit plots look wrong, Hogan notified
35 MeV bump in CSI energy/block plot, put note in log (XX.45)

=====================================================================
Follow up Notes on Previous Shift

It has been decided that we will run in BEAM_832 configuration even though
there is a known problem with E-W-U-D logic -- or rather a problem with the
V1 trigger sources used by that logic. V1 East is bad -- see JAG's notes
in log (XX.39) for more detailed explanation. It is not currently clear
whether the problem is caused in hardware or elsewhere. Work is in
progress.

========================================================================

Items to Follow up on Next Shift

Run in BEAM_832 config after Rick C. finishes with beam studies.

If module 2365 is swapped, return old module to PREP with form
currently attached to new module.